Cryptology Assignment

Due Friday, March 27, at 12:15.

Note that this is an actual assignment which must be approved in order to take the exam in June. You may not work with others on the assignment. It must be turned in by 12:15 on Friday, March 27. If it is late, it will not be accepted. There will be no chance to redo the assignment if it is not acceptable the first time.

Do problems 4.7, 4.8 and 4.9 in the textbook. For 4.7b, I recommend using the PowerMod function in Mathematica. For 4.8b, note that 945 does not work as an exponent. Try using b_{1} = 943 and y_{1} = 12921.